• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kuss

    Log entry time 09:23:49 on October 11,1999

    Entry number 28172

    keyword=target controls

    There is no connection to ITC502_1, the ITC502 which controls the three Cernox and the low power heater of Loop 1. IOC reboot didn't solve the problem, so it seems to be the hardware, not EPICS. Loop 1 is not in use, and those sensors are out-of-range anyway, so it's not critical. Will check tomorrow.